1. Suppose v,w element of V. Explain why there exists a unique x element of V such that v +3x = w.
1
Expert's answer
2021-05-12T03:43:35-0400

Suppose "v,w" element of "V". Let us show that there exists a unique "x" element of "V" such that "v +3x = w" . Let "x_1" and "x_2" are the solutions of the equation "v +3x = w". Then "v +3x_1 = w" and "v +3x_2 = w". It follows that "v +3x_1 =v +3x_2." Since "(V,+)" is a group, using left cancellation law we conclude that "3x_1 =3x_2". After multiplying by "\frac{1}{3}" from left we have that "\frac{1}{3}(3x_1) =\frac{1}{3}(3x_2)."

Then using compatibility of scalar multiplication with field multiplication we conclude that "(\frac{1}{3}3)x_1 =(\frac{1}{3}3)x_2," and hence "1\cdot x_1=1\cdot x_2." Using property of identity element of scalar multiplication we conclude that "x_1=x_2." Consequently, there exists a unique "x" element of "V" such that "v +3x = w" .